OpenStudy (anonymous):

PLEASE HELP, I HAVE A QUIZ IN 30 MINS Write the equation of a sine function in standard form with the following characteristics: Max: 39 Min: 29 Period: 24 Horizontal Shift: 6 units to the right Vertical reflection

OpenStudy (gorv):

we write function of sine as

OpenStudy (gorv):

A*sinBx

OpenStudy (gorv):

A=amplitude

OpenStudy (gorv):

sorry A*sin(Bx+C)

OpenStudy (gorv):

period= 2pi*B shift=C

OpenStudy (gorv):

shift to right means c=-6

OpenStudy (gorv):

A*sin(Bx-6)

OpenStudy (gorv):

period =2*pi/B=24

OpenStudy (gorv):

\[\frac{ 2* \pi }{ B}= 24\]

OpenStudy (gorv):

\[B=\frac{ 2*\pi }{ 24 }=\frac{ \pi }{ 12 }\]

OpenStudy (gorv):

\[A*\sin(\frac{ \pi*x }{ 24 }-6)\]

OpenStudy (gorv):

and reflection along y axis means multiply x by negative sign

OpenStudy (gorv):

amplitude = (max-min)/2

OpenStudy (gorv):

A=(39-29)/2=5 so your function \[-5\sin(\frac{ \pi x }{ 12}-6)\]

OpenStudy (gorv):

@maysa13

OpenStudy (anonymous):

Thank you so much gorv!!!